# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v2

@@ 3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
We use Monte Carlo tree search (MCTS) to solve the MDP of directed evolution. To effectively select actions, some MCTS works (rosin_multi-armed_2011; silver_mastering_2016) train neural networks based on expert data for experienced guidance. Similarly, to enable efficient guidance of protein mutation, AlphaDE employs a fine-tuned protein language model to predict the next mutation action (the residue position and residue type) given the current protein sequence. At each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type on each residue position as guidance for MCTS.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate a protein sequence in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (romero_exploring_2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for mutation guidance. Next, we introduce the MCTS part of AlphaDE in the background of protein directed evolution.
@@
  • Selection. Each iteration starts from sequence $s_{\tau}$, recursively selecting the best child node until reaching a leaf node $a_{\tau+l}$, i.e., a node that has not been expanded or terminated, after $l$ selections. At each selection step $t\in[1,l]$, a selection criterion determines the best child node to be chosen, which balances exploitation and exploration to avoid being trapped in local optimums. We use Predictor with Upper Confidence bounds applied to Trees (PUCT) (rosin_multi-armed_2011) as the selection criterion for each candidate child node as
@@
    where the constant $c$ controls the exploration degree. $W_{a}$ is the cumulative reward of node $a$, NN is the total visit count, and $N_{a}$ is the visit count for node $a$. The term $\frac{\sqrt{N}}{1+N_{a}}$ encourages selection of less-visited nodes, while $P_{plm}(a|s_{\tau+t-1})$ favors beneficial mutations predicted by the fine-tuned protein language model. Furthermore, $\frac{W_{a}}{N_{a}}$ promotes exploitation of the nodes with high protein function fitness.
@@
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination, then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (silver_general_2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network.
@@
    where $N_{a}$ is the visit count and $W_{a}$ is the cumulative reward of node $a$. For each selection $t\in[1,l]$, the statistics of node $a_{\tau+t}$ are updated by adding the rollout reward of $a_{\tau+l}$ to $W_{a}$ and increasing $N_{a}$ by 1.
@@ 4 Experiments
In this section, we conduct computational experiments to validate the effectiveness of AlphaDE. First, in Section 4.1, we benchmark AlphaDE with various directed evolution algorithms on extensive tasks. Second, we study AlphaDE under different settings of homology sequence availability, such as few-shot fine-tuning in Section 4.2 and zero-shot scaling with model sizes in Section 4.3. Finally, we utilize AlphaDE to computationally condense the protein sequence space of the fluorescent protein in Section 4.4. The ablation study of AlphaDE is provided in Appendix F.
@@ 4.1 Benchmark experiments
We compare AlphaDE with various baselines including LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, and EvoPlay for running. We take the results of PEX (one version with CNN model and another with MuFacNet model), AdaLead, CbAS, DbAS, and DyNA-PPO from Ren et al. (ren_proximal_2022) as we use the same benchmark setting, such as the starting sequence and fitness oracle. We run BO and CMA-ES with the implementations in FLEXS (sinai_adalead_2020). We average the results of AlphaDE, LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, EvoPlay, BO, and CMA-ES over 5 independent trials. For all methods, we set the same oracle budget of 1000 queries. For AlphaDE, here we use the pretrained ESM2-series models (lin_evolutionary-scale_2023) for fine-tuning as it is one of the most powerful and widely used protein language models with various model sizes for study convenience. Specifically, in the standard AlphaDE, we use ESM2-35M for fine-tuning as its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence for optimization is the sequence with the lowest protein fitness value, which is the same as the starting sequences in TAPE oracles. The benchmark results for these eight tasks are summarized in Table 1.
Table 1: Benchmarking in-silicon directed evolution methods. We present the maximum fitness in 1000 black-box oracle queries.
@@
AlphaDE 3.86 7.95 1.22 7.75 0.24 0.04 1.47 2.97
LatentDE (tran_latentde_2025) 3.79 4.72 1.20 3.88 -3.34 -1.13 0.58 1.29
MLDE (tran_protein_2025) 2.04 -3.38 0.08 3.84 -1.38 -0.19 0.92 2.75
TreeNeuralTS (qiu_tree_2024) 2.44 2.47 0.27 0.79 -0.22 0.04 1.02 2.98
TreeNeuralUCB (qiu_tree_2024) 2.37 3.85 0.19 0.70 -0.19 0.04 1.27 2.95
PEX (MuFacNet) (ren_proximal_2022) 3.12 4.45 0.27 2.22 0.16 0.04 1.23 2.97
PEX (CNN) (ren_proximal_2022) 2.97 2.52 0.19 2.21 -0.11 0.03 1.27 2.97
EvoPlay (wang_self-play_2023) 1.72 -3.45 0.01 -0.40 -0.88 -1.09 0.34 1.87
EvoPlay (wang_self-play_2023) 1.72 -3.45 0.01 -0.40 -0.88 -1.09 0.34 1.87
AdaLead (sinai_adalead_2020) 2.61 -2.33 0.09 0.16 -0.86 -0.72 1.09 2.91
DyNA-PPO (angermueller_model-based_2020) 1.84 -3.22 0.03 -0.20 -2.13 -0.32 0.42 2.17
DbAS (brookes_design_2018) 2.30 -2.43 0.10 0.23 -2.30 -0.35 0.90 2.85
CbAS (brookes_conditioning_2019) 2.22 -2.50 0.10 0.19 -2.26 -0.22 0.92 2.87
BO (snoek_practical_2012) 1.57 -3.84 0.02 -0.33 -5.71 -1.45 0.39 0.02
CMA-ES (hansen_completely_2001) 1.60 -3.50 0.02 -0.09 -7.92 -1.29 0.53 -0.01
The results in Table 1 show that AlphaDE significantly outperforms various baselines in most tasks. For example, in the task of AAV, the second best method MLDE obtains a fitness value of 4.72 while AlphaDE achieves 7.95 with an improvement of 68.43%. Similar situations also happen in the task of E4B and AMIE, which exhibit AlphaDE’s great evolution ability by combining effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model). At the same time, we also examine the diversity of evolved sequences by AlphaDE in Appendix I. AlphaDE with different ESM2 model sizes is tested in Appendix D. The ablation study of AlphaDE on the fine-tuning step and MCTS inference step is given in Appendix F. Benchmark results using another oracle based on ESM-1b further validate AlphaDE as shown in Appendix G. Moreover, the study of hyperparameter sensitivity of AlphaDE is available in Appendix H, while AlphaDE’s computation efficacy is provided in Appendix L.
At the same time, we also test AlphaDE with other protein language models such as ProtBert (elnaggar_prottrans_2022) (model size 420M) and ESM-1b (rives_biological_2021) (model size 650M), which are trained with the standard BERT (devlin_bert_2019) architecture using the UniRef100 (suzek_uniref_2007) database and with the optimized RoBERTa (liu_roberta_2019) architecture using the UniRef50 (suzek_uniref_2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. We see that AlphaDE with ProtBert and AlphaDE with ESM-1b perform differently in some tasks, suggesting their evolution ability varies in different protein families. Generally, AlphaDE with the two different protein language models achieves superior performance, showing AlphaDE’s compatibility. Meanwhile, the fine-tuning step is the key to boosting performance, as directly using the pretrained versions of protein language models leads to a lot of fitness loss.
Table 2: Results of AlphaDE with ProtBert and ESM-1b models. We present the maximum fitness scores obtained in 1000 black-box oracle queries. Results are averaged over five independent trials.
@@ 4.2 Few-Shot Fine-Tuning Experiments
One of the biggest advantages of large language models is that they are few-shot learners (brown_language_2020). Here we investigate whether this advantage is held in the protein language model for protein directed evolution tasks. We use the ESM2-35M model as the base model for few-shot fine-tuning and fine-tune it with different numbers of protein sequences including 16, 32, 64, 128, 256, 512, and 1024. These protein sequences for few-shot fine-tuning are randomly sampled from the whole dataset and are used to fine-tune ESM2-35M with 3 epochs. The results of AlphaDE with different few-shot fine-tuned ESM2-35M models are given in Figure 2. We see that, even with dozens of protein sequences, the evolution ability of AlphaDE could be greatly improved, and the performance increases with the number of fine-tuning protein sequences.
@@
Figure 3: AlphaDE with fine-tuned ESM2-35M, which are fine-tuned with different numbers of sequences randomly sampled from the bottom 20% data. The “Max Bottom 20%” value denotes the maximum fitness value in the bottom 20% data. 95% confidence intervals are shadowed.
Although AlphaDE does not utilize fitness labels of protein sequences in the pretrain, fine-tune, or MCTS inference, we additionally test whether the protein fitness influences the performance of AlphaDE by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3 and are very similar to Figure 2, showing that AlphaDE’s evolution capability is unaffected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly fitness measurement assays. We also show that AlphaDE works with the sequences from homology searching such as HMMER (potter_hmmer_2018) in Appendix J.
@@ 4.3 Zero-Shot Scaling Experiments on Model Size
Figure 4: AlphaDE’s performance scales with sizes of pretrained protein language models. The black dashed line indicates AlphaDE with fine-tuned ESM2-35M. 95% confidence intervals are shadowed.
Next, we study how AlphaDE scales with pretrained protein language model sizes in a zero-shot setting, testing ESM2 models from 8M to 15B. Results in Figure 4 show that, in most tasks except avGFP and AAV, AlphaDE’s performance roughly increases with pretrained model sizes. This is different from Table 3, where the fine-tuned model size has little effect. Interestingly, with the largest pretrained 15B model, AlphaDE’s performance matches with its fine-tuned ESM2-35M version on tasks of TEM, AMIE, and UBE2I. However, in most tasks, the pretrained models, even 15B, perform worse than the standard AlphaDE with fine-tuned ESM2-35M. We see two points here. First, larger pretrained models encode more evolutionary information, enhancing AlphaDE’s evolution ability in most cases. Second, fine-tuned small models with homologous sequences provide strong task-specific evolution ability, underscoring the necessity of fine-tuning.
@@ 4.4 Condensing Protein Sequence Space
Figure 5: The illustrated process of AlphaDE to condense the sequence space of avGFP. The evolution trajectory is sampled during one trial, and the first is the predicted structure of the starting sequence.
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (abramson_accurate_2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the comparison of AlphaDE without the fine-tuning step and the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ 5 Conclusion
In this work, we propose AlphaDE for the challenging protein directed evolution problem. AlphaDE consists of the fine-tuning step and the MCTS test-time inference step based on the protein language models to attain superior evolution ability. Benchmark experiments on eight tasks show that AlphaDE achieves state-of-the-art performance compared with various baselines, even in a few-shot setting. Additionally, we also demonstrate that AlphaDE could be utilized to computationally condense the protein sequence space.
For future work, on the one hand, integrating large natural language models into AlphaDE to provide the explainability of the evolution process is interesting. On the other hand, applying AlphaDE to industrial applications, such as improving the activity of enzymes, has great economic prospects.

### v3

@@ 3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
We use Monte Carlo tree search (MCTS) to solve the MDP of directed evolution with guidance from protein language models, inspired by MCTS-based test-time reasoning boosts in large language models that tree-based search prompts better solutions (zhou_language_2024; guan_rstar-math_2025). Diverging from prior MCTS works that train neural networks on expert data from scratch for action guidance (rosin_multi-armed_2011; silver_mastering_2016), our approach leverages a fine-tuned protein language model to provide evolutionary guidance for mutation decisions, which predicts the next mutation action (i.e., the residue position and residue type) given the current protein sequence. Specifically, at each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type at each residue position to guide the MCTS search.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate a protein sequence in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (romero_exploring_2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for mutation guidance. Next, we explain the MCTS part of AlphaDE in the background of protein directed evolution.
@@
  • Selection. Each iteration starts from sequence $s_{\tau}$, recursively selecting the best child node until reaching a leaf node $a_{\tau+l}$, i.e., a node that has not been expanded or terminated, after $l$ selections. At each selection step $t\in[1,l]$, a selection criterion determines the best child node to be chosen, which balances exploitation and exploration to avoid being trapped in local optimum. We use Predictor with Upper Confidence bounds applied to Trees (PUCT) (rosin_multi-armed_2011) as the selection criterion for each candidate child node as
@@
    where a constant $c$ controls the exploration degree. $W_{a}$ is the cumulative reward of node $a$, NN is the total visit count, and $N_{a}$ is the visit count for node $a$. The term $\frac{\sqrt{N}}{1+N_{a}}$ encourages selection of less-visited nodes, while $P_{plm}(a|s_{\tau+t-1})$ favors beneficial mutations indicated by the fine-tuned protein language model. On the other hand, $\frac{W_{a}}{N_{a}}$ encourages exploitation of the nodes with high average protein fitness.
@@
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination at step $T$. Then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (silver_general_2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network.
@@
    where $N_{a}$ is the visit count and $W_{a}$ is the cumulative reward of node $a$. For each selection $t\in[1,l]$, the statistics of node $a_{\tau+t}$ are updated by adding the rollout reward of $a_{\tau+l}$ to $W_{a}$ and by increasing $N_{a}$ by 1.
@@ 4 Experiments
In this section, we conduct comprehensive experiments to validate the effectiveness of AlphaDE. First, in Section 4.1, we benchmark AlphaDE with various directed evolution algorithms on extensive tasks. Second, we study AlphaDE under different settings of homology sequence availability, such as few-shot fine-tuning in Section 4.2 and zero-shot scaling with model sizes in Section 4.3. Finally, we utilize AlphaDE to computationally condense the protein sequence space of the fluorescent protein in Section 4.4. The ablation study of AlphaDE is provided in Appendix F.
@@ 4.1 Benchmark experiments
We compare AlphaDE with various baselines including LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX (one version with a CNN model and another with a MuFacNet model), and EvoPlay for running. We run AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES with the implementations in FLEXS (sinai_adalead_2020), following the same benchmark setting, such as the starting sequence and fitness oracle as Ren et al. (ren_proximal_2022). We average the results of AlphaDE and all baselines over 5 independent trials. For all methods, we set the same oracle budget of 1000 queries. For AlphaDE, here we use the pretrained ESM2-series models (lin_evolutionary-scale_2023) for fine-tuning due to its strong performance, wide adoption, and availability in various model sizes, which is convenient for our study. Specifically, in the standard AlphaDE, we use ESM2-35M for fine-tuning due to its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence to evolve is the sequence with the lowest protein fitness. The benchmark results for eight tasks are summarized in Table 1, which includes mean values and standard deviations.
Table 1: Benchmarking in-silicon directed evolution methods. We present the maximum fitness obtained in 1000 black-box oracle queries. Results are presented as mean with standard deviation over five independent trials.
@@
[…]
The results in Table 1 show that AlphaDE outperforms various baselines in most tasks. For example, in the task of AAV, the second-best method, MLDE, obtains a fitness value of 4.72 while AlphaDE achieves 7.95 with an improvement of 68.43%. A similar situation also happens in the task of E4B, which exhibits AlphaDE’s great evolution ability by combining effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model). At the same time, we also examine the diversity of evolved sequences by AlphaDE in Appendix I, which shows nearly 100% sequence diversity across running trials. AlphaDE with different ESM2 model sizes is tested in Appendix D. Meanwhile, the ablation study of AlphaDE on the fine-tuning step and MCTS inference step is given in Appendix F. Benchmark results using another oracle based on ESM-1b further validate AlphaDE as shown in Appendix G. Moreover, the study of hyperparameter sensitivity of AlphaDE is available in Appendix H, while AlphaDE’s computation efficacy is provided in Appendix L.
We also test AlphaDE’s compatibility with other protein language models such as ProtBert (elnaggar_prottrans_2022) (model size 420M) and ESM-1b (rives_biological_2021) (model size 650M), which are trained with the standard BERT (devlin_bert_2019) architecture using the UniRef100 (suzek_uniref_2007) database and with the optimized RoBERTa (liu_roberta_2019) architecture using the UniRef50 (suzek_uniref_2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. Generally, AlphaDE with the two different protein language models achieves superior performance, demonstrating AlphaDE’s compatibility. On the other hand, AlphaDE with ProtBert and ESM-1b perform differently in some tasks, suggesting the evolution ability of protein language models varies in different protein families. Meanwhile, the fine-tuning step is the key to boosting performance, as directly using the pretrained versions of models leads to huge fitness loss.
Table 2: Results of AlphaDE with ProtBert and ESM-1b models. We present the maximum fitness obtained in 1000 black-box oracle queries. Results are averaged over five independent trials.
@@ 4.2 Few-Shot Fine-Tuning Experiments
One of the biggest advantages of large language models is that they are few-shot learners (brown_language_2020). Here we investigate whether this advantage is held in the protein language model for directed evolution tasks. We use the ESM2-35M model as the base model for few-shot fine-tuning and fine-tune it with different numbers of protein sequences including 16, 32, 64, 128, 256, 512, and 1024. These protein sequences for few-shot fine-tuning are randomly sampled from the whole dataset and are used to fine-tune ESM2-35M with 3 epochs. The results of AlphaDE with different fine-tuned ESM2-35M models of different sequence numbers are given in Figure 2. We see that, even with dozens of protein sequences, the evolution ability of AlphaDE could be greatly improved, and the performance increases with the number of fine-tuning protein sequences.
@@
Figure 3: AlphaDE with fine-tuned ESM2-35M models of different numbers of sequences randomly sampled from the bottom 20% data. The “Max Bottom 20%” denotes the maximum fitness value in the bottom 20% data. 95% confidence intervals are shadowed.
Although AlphaDE does not utilize fitness labels of protein sequences in the fine-tuning and MCTS inference steps, we additionally test whether the protein fitness influences AlphaDE’s performance by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3. The curves are very similar to Figure 2, showing that AlphaDE’s evolution capability is not affected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly wet measurement assays.
We also show that AlphaDE works with the sequences from homology searching such as HMMER (potter_hmmer_2018) in Appendix J, where homology sequences are retrieved from a database.
@@ 4.3 Zero-Shot Scaling Experiments on Model Size
Figure 4: AlphaDE’s performance scales with the sizes of pretrained protein language models. The black horizontal dashed line indicates AlphaDE with fine-tuned ESM2-35M. 95% confidence intervals are shadowed.
Next, we study how AlphaDE scales with pretrained protein language model sizes in a zero-shot setting, testing ESM2 models from 8M to 15B. Results in Figure 4 show that, in most tasks except avGFP and AAV, AlphaDE’s performance roughly increases with pretrained model sizes. This is different from Table 3 in Appendix D, where the fine-tuned model size has little effect. Interestingly, with the largest pretrained 15B model, AlphaDE’s performance matches its fine-tuned ESM2-35M version on tasks of TEM, AMIE, and UBE2I. However, in most tasks, the pretrained models, even 15B, perform worse than the standard AlphaDE with fine-tuned ESM2-35M. We see two points here. First, larger pretrained models encode more evolutionary information, enhancing AlphaDE’s evolution ability in most cases. Second, the fine-tuned small models with homologous sequences provide strong task-specific evolution ability, underscoring the necessity of fine-tuning in AlphaDE’s pipeline.
@@ 4.4 Condensing Protein Sequence Space
Figure 5: The illustrated process of AlphaDE to condense the sequence space of avGFP. The evolution trajectory is sampled during one trial, and the first structure is the prediction of the starting sequence by AlphaFold 3.
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (abramson_accurate_2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the comparison to AlphaDE without the fine-tuning step and the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ 5 Conclusion
In this work, we propose AlphaDE to boost the in-silicon protein directed evolution. Following the paradigm of natural large language models, AlphaDE consists of a fine-tuning step and an MCTS test-time inference step based on the protein language models to attain superior evolution ability. Specifically, AlphaDE conducts MCTS to efficiently evolve proteins via residue mutations guided by the fine-tuned protein language model. Benchmark experiments on eight tasks show that AlphaDE achieves state-of-the-art performance compared with various baselines, even in the few-shot setting. Additionally, we demonstrate that AlphaDE supports computationally condensing the protein sequence space.
For future work, on the one hand, integrating large natural language models into AlphaDE to provide the explainability of the evolution process is interesting. On the other hand, applying AlphaDE to industrial applications, such as improving the activity of enzymes, has great economic prospects. Third, extending AlphaDE to the setting of multi-objective directed evolution is also a promising direction.